The following information is provided for Kelly Plumbing Supply Cash received from customers during Decenber 2017 Cash paid to suppliers for inventory during December 2017 . $387,000 131,000 Cash received from customers includes all $139,000 of the accounts receivable that were outstanding at November 30, 2017 Accounts recelvable at December 31, 2017 totaled $141.000. Accounts payable (to suppliers of inventory) decreased by $19,000 from November 30, 2017 to December 31, 2017. The balance in the inventory account decreased by $39,000 over the same period. Required What is gross profit for the month of December under accrual accounting?
On September 1,2017, Revsine Co. approved a plan to dispose of a segment of its business. Revsine expected that the sale would occur on March 31, 2018, at an estimated gain of $375,000. The segment had actual and estimated operating profits (losses) as ollows: Realized loss from 1/1/17 to 8/31/17 Realized loss from 9/1/17 to 12/31/17 Expected profit from 1/1/18 to 3/30/18 (300,000) (200,000) 400,000 The expected profit from 1/1/18 to 3/30/18 was based on Revsines expectations as of 12/31nz Assume the marginal tax rate is 30%. Required: In its 2017 income statement, what should Revsine report as profit or loss from discontinued operations (net of tax effects)?
Answer #1

1. Cash received from customers in Dec 2017 (a) = 387000.

Accounts receivable as on Nov 30th 2017 (b) = 139000

Cash Received for services rendered in Dec 2017(c) = (a-b) = 248000.

  Accounts receivable as on Dec 30th 2017 (d) = 141000.

  Value of Services rendered In Dec 2017 = (c+d) = $ 389000 (under accrual basis).

Balance in Inventory Account decreased by 39000, That means we have utilised inventory worth $ 39000 in Dec 2017.

Gross Profit = 389000-39000 =$ 350000.

The fact that Accounts payable have reduced by $ 19000 will not have any bearing on Gross Profit as it a reduction in liability, therefore an appropriation of cash received towards payables, but not a revenue or expense.

2. As Revsine Co have formalized a plan to dispose of a Segment of its business by the reporting date, the same should be reported in the financial statement as "Discontinued Operations"

Total realized Loss in the Year 2017 = $ 300000 + $ 200000 = $ 500000.

Tax Rate = 30%.

Loss to be reported under Discontinued operations = 500000* (1-0.30) = $ 350,000\-

The fact that the company is expected to gain on sale on 31st March 2018 and is expecting a gain on sale will have no bearing on the reporting for the year 2017.

However the same can be informed to the stakeholders in an appropriate way.